# CFT Constraints on Parity-odd Interactions with Axions and Dilatons

### v1

@@ Abstract
We illustrate how the Conformal Ward Identities (CWIs) in momentum space completely determine the structure of a parity-odd 3-point correlator involving currents, energy momentum tensors and at least one scalar operator in $d=4$. Conformal invariance fixes almost all such possible correlators to vanish. The only exceptions are given by the $\langle JJO\rangle_{odd}$ and the $\langle TTO\rangle_{odd}$ which in momentum-space are protected by chiral and conformal anomalies. Specifically, one can obtain a non-vanishing solution by considering scalar operators such as $O=\nabla\cdot J_{A}$, $O=g_{\mu\nu}T^{\mu\nu}$ or their shadow transforms. We comment on the implications of these results that constrain the coupling of axions and dilatons in a conformal phase of the early universe.
@@ 1 Introduction
Determining correlation functions in momentum space is, however, significantly more intricate than in position space. The use of conformal Ward identities to determine the structure of three-point functions in momentum space was independently introduced in [1] and [2], with the latter extending the methodology to tensor cases. Renormalized parity even correlators, type-A and type-B Weyl anomalies have also been extensively discussed [3, 4, 5] [6]. Perturbative analysis of some important correlators such as the $\langle TJJ\rangle$, in the parity-even sector had been investigated in the conformal and non-conformal limits in previous works in QED and QCD [7, 8, 9].
Recently, parity-odd correlators and non-conserved currents have also been the focus of attention [10, 11, 12, 13]. In particular, in [14, 15, 16] parity-odd correlators have been explored in relation to chiral and conformal anomalies. The latter papers demonstrate how such anomalies are responsible for generating non-vanishing parity-odd correlators. In the absence of chiral anomalies, correlators like $\langle J_{V}J_{V}J_{A}\rangle$, $\langle J_{A}J_{A}J_{A}\rangle$, $\langle TTJ_{A}\rangle$ where $J_{V}$ and $J_{A}$ are vector and axial-vector current respectively and $T$ the stress energy tensor - would entirely vanish in a CFT. Likewise, without parity-odd conformal anomalies, the $\langle TJJ\rangle_{odd}$ correlator, where $J$ is a generic current of any parity, is zero.
@@
In this paper we further examine the conformal constraints for parity-odd correlators in momentum space in $d=4$. We focus in particular on all the possible 3-point correlators built with currents, energy momentum tensors and at least one scalar operator.
@@
The nature of the pseudoscalar operator introdued in our investigation is general, and can be of any scaling dimension $\Delta$. The most interesting cases, from our perspective, are for $\Delta=4$, where $O\sim\partial J_{5}$ and $O\sim T^{\mu}_{\mu}$, which carry direct physical implications.
@@
In the case of the $\langle JJO\rangle$, for example, our results remain identical even if $O$ and $J$ are both of odd parity, as the behavior of the correlator under parity is a global property of the correlator and not necessarily of a single operator in the triple. As we will show, the solution of the conformal and parity constraints naturally requires that their tensor structure is essentially that induced by a chiral anomaly. A modification of this behavior is encountered in the shadow solutions of the CWIs, where a primary operator $O$ of scaling dimension $\Delta$ is replaced by its shadow-transformed expression of dimension $d-\Delta$. The corresponding operators, in this case, can be defined by a nonlocal shadow transform.
@@
Our work is organized as follows. In Section 2, we briefly discuss correlators with two and three scalar operators. In Section 3, we then derive the solution of the CWIs for the $\langle TJO\rangle$. This correlator interpolates with the (parity-odd) amplitude for an axion/dilaton particle transitioning into a spin-1 state in the presence of an off-shell external gravitational field. As we will demonstrate, this amplitude vanishes identically in a conformal phase of the early universe. Then, in Section 4, we turn to reinvestigate the $\langle JJO\rangle$ and the $\langle TTO\rangle$ correlator. For the latter, we show that the CWIs admit additional solutions that had not been identified in a previous analysis [15]. Such solutions are connected by shadow transforms of the primary scalars. Lastly, before presenting our conclusions, in Section 5, we will discuss the relevance of this study at the cosmological level, specifically regarding the analysis of chiral and dilatation current anomalies in the early universe.
@@ 2 Three-point functions with multiple scalars
We start by considering correlators with at least two scalar operators $O(p_{i})$. Before examining the conformal constraints, we write the most general form of the correlators in terms of form factors and tensorial structures. In this respect, the vanishing of such correlators in the parity-odd sector is rather straightforward, due to symmetry. Indeed, since we are dealing with four-dimensional parity-odd correlators, the tensorial structures need to include an $\varepsilon^{\alpha_{1}\alpha_{2}\alpha_{3}\alpha_{4}}$ tensor. It is to figure out that this condition cannot be satisfied in this case, thereby giving the vanishing relations
@@ 3 The conformal $\langle TJO\rangle_{odd}$
in order to sperate its several (longitudinal, trace and transverse-traceless) sectors, expanded with respect to the external momenta. Because of eq. (2), the correlator is composed only of a transverse-traceless part which can be written as
@@ 3.1 Dilatations and special conformal equations
where $C_{i}$ are scalar function that depend on the form factor $A$ and its derivatives with respect to the momenta. The tensor structures listed in the equation above are not all independent and can be simplified in order to find a minimal decomposition, using the following Schouten identities
@@
00 0 $=\epsilon^{[p_{1}p_{2}\mu_{1}\mu_{2}}\delta^{\kappa]\nu_{1}}.$
@@
where we have introduced the operator
@@
These operators [2] allow to relate the differential content of the equations for the form factors, which are combinations of Appell hypergeometric functions [1, 32] of two variables, briefly discussed in Appendix A.
@@ 3.2 Solving the CWIs
The solution of the conformal Ward identities for the $\langle TJO\rangle$ can be written in terms of integrals involving a product of three Bessel functions, namely 3K integrals [2, 33], as illustrated Appendix A. 4K extensions of these formulations appear in the large energy $s$ and momentum transfer $t$ limits of four-point functions in CFT, as solutions special combinations of Lauricella functions [34]. In the 3K case, we recall the definition
@@
where $K_{\nu}$ is a modified Bessel function of the second kind
@@
We will also use the reduced version of the 3K integral defined as
@@
The 3K integrals satisfy an equation analogous to the dilatation equation with scaling degree [2]
@@
From this analysis, it is simple to relate the form factor $A$ to the 3K integrals. Indeed, the dilatation WI (9) tells us that the form factor $A$ can to be written as a combination of integrals of the following type
@@
For a more detailed review of the topic, see appendix A and [2, 3, 33]. If the above condition is satisfied, we need to regularize the integral. This can be done by shifting the parameters of the 3K integrals as
@@
In general, the regularization parameters $u$ and $v_{i}$ are arbitrary. However, in certain specific cases, there can be some constraints on them.
@@ 4 The conformal $\langle JJO\rangle_{odd}$ and $\langle TTO\rangle_{odd}$  : a reappraisal
Let’s now discuss the $\langle JJO\rangle_{odd}$ and $\langle TTO\rangle_{odd}$ correlators, which have been investigated in [15]. The procedure follows the same approach of the $\langle TJO\rangle_{odd}$ discussed in the previous section.
First, we will briefly review the analysis of the $\langle JJO\rangle_{odd}$ correlator. Then, we will proceed with a re-evaluation of the $\langle TTO\rangle_{odd}$, as the primary and secondary equations in the published version of [15] were not correctly identified. Here, we present the correct version of these equations and solve them accordingly. The previous solution reported in [15] for the $\langle TTO\rangle_{odd}$ with $\Delta_{3}=4$ remains valid, and the conclusions and interpretations of that paper are unaffected. However, these corrections have led us to discover new non-vanishing solutions for the $\langle TTO\rangle_{odd}$ characterised by different values of $\Delta_{3}$. Lastly, we further elaborate on all the solutions of both the $\langle JJO\rangle_{odd}$ and $\langle TTO\rangle_{odd}$, offering a new interpretation using the shadow transform.
@@ 4.2 The $\langle TTO\rangle_{odd}$
The most general solution of these equations can then be written in terms of the following combinations of 3K integrals
@@ 4.3 Connections with the anomalies
in the flat limit. By Fourier transforming such expressions, one obtains exactly the solutions in eq. (39) and (4.2) with $\Delta_{3}=4$. Note that $O=\nabla\cdot J_{A}$ is a primary operator since, by acting on it with the special conformal operator $\mathcal{K}$, we obtain a vanishing result
@@
Besides $O=\nabla\cdot J_{A}$, one could alternatively consider $O=g_{\mu\nu}T^{\mu\nu}$ as scalar operator with $\Delta_{3}=4$. The correlators would then be similarly protected by the possible existence of parity-odd trace anomalies in CFT
@@ 4.4 The shadow transforms and the $\Delta_{3}\leq 0$ solutions
The general solution of these correlators with $\Delta_{3}=4$ and $\Delta_{3}=0$ differ by a factor $p_{3}^{4}$ Although the condition $\Delta_{3}=0$ is non-physical and violate unitarity, we can now ask ourself if there are any justifications for the existance of such non-vanishing solutions in CFT. As we will see, the conformal and chiral anomalies can still be responsible for such solutions. To demonstrate this, we need to introduce the concept of shadow transform, which we will briefly review [35, 36, 37].
@@
Given this definition, we now reexamine the solutions of the $\langle JJO\rangle_{odd}$ and $\langle TTO\rangle_{odd}$ in eq. (39) and (4.2). For both correlators, the solution with $\Delta_{3}=0$ is connected to the one with $\Delta_{3}=4$ through a shadow transformation of the scalar operator $O$.

### v2

@@ Abstract
We illustrate how the conformal Ward identities (CWIs) in momentum space completely determine the structure of a parity-odd 3-point correlator involving currents, energy-momentum tensors and at least one scalar operator in $d=4$. Conformal invariance fixes almost all such possible correlators to vanish. The only exceptions are given by the $\langle JJO\rangle_{odd}$ and the $\langle TTO\rangle_{odd}$ which in momentum space are protected by chiral and conformal anomalies. Specifically, one can obtain a non-vanishing solution by considering scalar operators such as $O=\nabla\cdot J_{A}$, $O=g_{\mu\nu}T^{\mu\nu}$ or their shadow transforms. We comment on the implications of these results that constrain the coupling of axions and dilatons in a conformal phase of the early universe.
@@ 1 Introduction
Determining correlation functions in momentum space is, however, significantly more intricate than in position space. The use of conformal Ward identities to determine the structure of three-point functions in momentum space was independently introduced in [1] and [2], with the latter extending the methodology to tensor cases. Renormalized parity even correlators, type-A and type-B Weyl anomalies have also been extensively discussed [3, 4, 5, 6]. Furthermore, previous perturbative analyses in QED and QCD have examined important correlators such as the parity-even $\langle TJJ\rangle$, involving the energy-momentum tensor $T$ and currents $J$, in both conformal and non-conformal settings [7, 8, 9].
Recently, parity-odd correlators and non-conserved currents have also been the focus of attention [10, 11, 12, 13]. In particular, in [14, 15, 16] parity-odd correlators have been explored in relation to chiral and conformal anomalies. The latter papers demonstrate how such anomalies are responsible for generating non-vanishing parity-odd correlators. In the absence of chiral anomalies, correlators like $\langle J_{V}J_{V}J_{A}\rangle$, $\langle J_{A}J_{A}J_{A}\rangle$, $\langle TTJ_{A}\rangle$ — where $J_{V}$ and $J_{A}$ are vector and axial-vector currents respectively — would entirely vanish in a CFT. Likewise, without parity-odd conformal anomalies, the $\langle TJJ\rangle_{odd}$ correlator, where $J$ is a generic current of any parity, is zero.
@@
In this paper we further examine the conformal constraints for parity-odd correlators in momentum space in $d=4$. We focus in particular on all the possible 3-point correlators built with currents, energy-momentum tensors and at least one scalar operator.
@@
The nature of the pseudoscalar operator introduced in our investigation is general, and can be of any scaling dimension $\Delta$. The most interesting cases, from our perspective, are for $\Delta=4$, where $O\sim\partial J_{5}$ and $O\sim T^{\mu}_{\mu}$, which carry direct physical implications.
@@
In the case of the $\langle JJO\rangle_{odd}$, for example, our results remain identical even if $O$ and $J$ are both of odd parity, as the behavior of the correlator under parity is a global property of the correlator and not necessarily of a single operator in the triple. As we will show, the solution of the conformal and parity constraints naturally requires that their tensor structure is essentially that induced by a chiral anomaly. A modification of this behavior is encountered in the shadow solutions of the CWIs, where a primary operator $O$ of scaling dimension $\Delta$ is replaced by its shadow-transformed expression of dimension $d-\Delta$. The corresponding operators, in this case, can be defined by a non-local shadow transform.
@@
Our work is organized as follows. In Section 2, we describe the CWIs in coordinate and in momentum space. In Section 3, we briefly discuss parity-odd correlators with two and three scalar operators. In Section 4, we then derive the solution of the CWIs for the $\langle TJO\rangle_{odd}$. This correlator interpolates with the (parity-odd) amplitude for an axion/dilaton particle transitioning into a spin-1 state in the presence of an off-shell external gravitational field. As we will demonstrate, this amplitude vanishes identically in a conformal phase of the early universe. Then, in Section 5, we turn to reinvestigate the $\langle JJO\rangle_{odd}$ and the $\langle TTO\rangle_{odd}$ correlator. For the latter, we show that the CWIs admit additional solutions that had not been identified in a previous analysis [15]. Such solutions are connected by shadow transforms of the primary scalars. Lastly, before presenting our conclusions, in Section 6, we discuss the relevance of this study at the cosmological level, specifically regarding the analysis of chiral and dilatation current anomalies in the early universe.
2 Conformal Ward identities from coordinate to momentum space
The Poincaré group is a fundamental symmetry group in physics, encompassing translations and Lorentz transformations. These transformations leave the metric invariant, making the Poincaré group essential in the study of special relativity and quantum field theory. However, certain physical theories, such as those involving massless particles (like photons) or in the study of critical phenomena, exhibit a broader set of symmetries. This broader symmetry group is known as the conformal group, which extends the Poincaré group to include additional transformations that preserve the angle between spacetime intervals, rather than their exact length. In addition to the Poincaré transformations, the conformal group includes dilations and special conformal transformations. This extension is crucial for understanding systems with scale invariance, such as massless field theories, critical phenomena, and the geometry of spacetime in the context of the AdS/CFT correspondence.
Correlators that are invariant under the full conformal group need to satisfy peculiar relations that highly constrain their structure. Let us consider a scalar $n$-point correlator function $\left\langle{O}_{1}\left(x_{1}\right)\ldots{O}_{n}\left(x_{n}\right)\right\rangle$ where $\Delta_{1},\dots,\Delta_{n}$ represent the dimensions of the operators. The Ward identity for an infinitesimal transformation $g$ can be expressed as
$0=\sum_{j=1}^{n}\braket{{O}_{1}(x_{1})\dots\delta_{g}{O}_{j}(x_{j})\dots O_{n}(x_{n})},$ (1)
We now recall the conformal transformations rules
$\begin{split}[P_{\mu},{O}_{j}(x)]&=\partial_{\mu}{O}_{j}(x),\\ [D,{O}_{j}(x)]&=(\Delta_{j}+x^{\mu}\partial_{\mu}){O}_{j}(x),\\ [L_{\mu\nu},{O}_{j}(x)]&=\left(\mathcal{S}_{\mu\nu}+x_{\nu}\partial_{\mu}-x_{\mu}\partial_{\nu}\right){O}_{j}(x),\\ [K_{\mu},{O}_{j}(x)]&=\left[2x_{\mu}(x\cdot\partial)-x^{2}\partial_{\mu}+2\Delta_{j}x_{\mu}-2x^{\nu}\mathcal{S}_{\mu\nu}\right]{O}_{j}(x).\end{split}$ (2)
Here, $\mathcal{S}_{\mu\nu}$ is a finite-dimensional representation matrix of rotations, which determines the spin of the field and vanishes in the case of a scalar operator. The dilatation Ward identity can then be expressed as
$0=\left[\sum_{j=1}^{n}\Delta_{j}+\sum_{j=1}^{n}x_{j}^{\alpha}\frac{\partial}{\partial x_{j}^{\alpha}}\right]\left\langle{O}_{1}\left(x_{1}\right)\ldots{O}_{n}\left(x_{n}\right)\right\rangle$ (3)
while the special conformal Ward identity is given by
$0=\sum_{j=1}^{n}\left[2\Delta_{j}x_{j}^{k}+2x_{j}^{\kappa}x_{j}^{\alpha}\frac{\partial}{\partial x_{j\,\alpha}}-x_{j}^{2}\frac{\partial}{\partial x_{jk}}\right]\left\langle{O}_{1}\left(x_{1}\right)\ldots{O}_{n}\left(x_{n}\right)\right\rangle$ (4)
where $k$ is a free Lorentz index. In the case of tensor operators one needs to include an additional term to the previous equation, i.e. the contribution $\mathcal{S}_{\mu\nu}$ in eq. (2). For example, the special CWI of the $\langle TJO\rangle$ correlator takes the following form
$0=$ $\sum_{j=1}^{3}\left[2\Delta_{j}x_{j}^{k}+2x_{j}^{\kappa}x_{j}^{\alpha}\frac{\partial}{\partial x_{j\,\alpha}}-x_{j}^{2}\frac{\partial}{\partial x_{jk}}\right]\left\langle T^{\mu_{1}\nu_{1}}\left(x_{1}\right)J^{\mu_{2}}\left(x_{2}\right)O\left(x_{3}\right)\right\rangle$ (5)
$+4\bigg{(}(x_{1})_{\alpha}\delta^{\kappa\left(\mu_{1}\right.}-\delta_{\alpha}^{\kappa}\,x_{1}^{\left(\mu_{1}\right.}\bigg{)}\langle T^{\left.\nu_{1}\right)\alpha}\left(x_{1}\right)J^{\mu_{2}}\left(x_{2}\right)O\left(x_{3}\right)\rangle$
$+2\bigg{(}(x_{2})_{\alpha}\delta^{\kappa\mu_{2}}-\delta_{\alpha}^{\kappa}\,x_{2}^{\mu_{2}}\bigg{)}\left\langle T^{\mu_{1}\nu_{1}}\left(x_{1}\right)J^{\alpha}\left(x_{2}\right)O\left(x_{3}\right)\right\rangle.$
The CWIs above may be Fourier transformed in a similar manner to that discussed in [32]. Due to the translation invariance the position space correlators depend only on the differences $x_{j}-x_{n}$. Therefore, we can set $x_{n}=0$ and take
${p}_{n}=-\sum_{j=1}^{n-1}{p}_{j}$ (6)
The dilatation equation in momentum space takes the form
$\left[\sum_{j=1}^{n}\Delta_{j}-(n-1)d-\sum_{j=1}^{n-1}p_{j}^{\alpha}\frac{\partial}{\partial p_{j}^{\alpha}}\right]\left\langle{O}_{1}\left(p_{1}\right)\ldots{O}_{n}\left(p_{n}\right)\right\rangle=0$ (7)
while the special conformal Ward identity is given by
$0=\sum_{j=1}^{n-1}\left(2\left(\Delta_{j}-d\right)\frac{\partial}{\partial p_{j}^{\kappa}}-2p_{j}^{\alpha}\frac{\partial}{\partial p_{j}^{\alpha}}\frac{\partial}{\partial p_{j}^{\kappa}}+\left(p_{j}\right)_{\kappa}\frac{\partial}{\partial p_{j}^{\alpha}}\frac{\partial}{\partial p_{j\alpha}}\right)\left\langle{O}_{1}\left(p_{1}\right)\ldots{O}_{n}\left(p_{n}\right)\right\rangle$ (8)
Once again, in the tensorial case we need to include an additional term on the right-hand side of the special conformal equation. For example, in the case of the $\langle TJO\rangle$, the Ward identity takes the following form
$0=$ $\sum_{j=1}^{2}\left(2\left(\Delta_{j}-d\right)\frac{\partial}{\partial p_{j\kappa}}-2p_{j}^{\alpha}\frac{\partial}{\partial p_{j}^{\alpha}}\frac{\partial}{\partial p_{j\kappa}}+\left(p_{j}\right)^{\kappa}\frac{\partial}{\partial p_{j}^{\alpha}}\frac{\partial}{\partial p_{j\alpha}}\right)\left\langle T^{\mu_{1}\nu_{1}}\left(p_{1}\right)J^{\mu_{2}}\left(p_{2}\right)O\left(p_{3}\right)\right\rangle$ (9)
In the following sections, we will explore how the CWIs outlined above constrain parity-odd 3-point correlators involving at least one scalar operator. We will begin by decomposing the correlator into a longitudinal-trace and a transverse-traceless sector. Subsequently, we will solve the conformal equations to determine the most general structure for a correlator in a CFT, following the methodology introduced in [2]. As we will demonstrate, the solution to the CWIs in momentum space can be expressed using a specific class of functions known as 3K integrals, which involve products of Bessel functions.
@@ 3 Three-point functions with multiple scalars
We start by considering correlators with at least two scalar operators $O(p_{i})$. Before examining the conformal constraints, we write the most general form of the correlators in terms of form factors and tensorial structures. In this respect, the vanishing of such correlators in the parity-odd sector is rather straightforward, due to symmetry. Indeed, since we are dealing with four-dimensional parity-odd correlators, the tensorial structures need to include an $\varepsilon^{\alpha_{1}\alpha_{2}\alpha_{3}\alpha_{4}}$ tensor. It is easy to figure out that this condition cannot be satisfied in this case, thereby giving the vanishing relations
@@ 4 The conformal $\langle TJO\rangle_{odd}$
in order to separate its several (longitudinal, trace and transverse-traceless) sectors, expanded with respect to the external momenta. Because of eq. (11), the correlator is composed only of a transverse-traceless part which can be written as
@@ 4.1 Dilatations and special conformal equations
where $C_{i}$ are scalar functions that depend on the form factor $A$ and its derivatives with respect to the momenta. The tensor structures listed in the equation above are not all independent and can be simplified in order to find a minimal decomposition, using the following Schouten identities
@@
00 0 $=\epsilon^{[p_{1}p_{2}\mu_{1}\mu_{2}}\delta^{\kappa]\nu_{1}},$
@@
where we have introduced the operator [2]
@@
The operators $K_{ij}$ allow to relate the differential content of the equations for the form factors, which are combinations of Appell hypergeometric functions [1, 32] of two variables, briefly discussed in Appendix A.
@@ 4.2 Solving the CWIs
The solution of the conformal Ward identities for the $\langle TJO\rangle$ can be written in terms of integrals involving a product of three Bessel functions, namely 3K integrals [2, 33], as illustrated in Appendix A. 4K extensions of these formulations appear in the large energy $s$ and momentum transfer $t$ limits of four-point functions in CFT, as special combinations of Lauricella functions [34]. In the 3K case, we recall the definition
@@
where $K_{\nu}$ is the modified Bessel function of the second kind
@@
We will also use a reduced version of the 3K integral defined as
@@
The 3K integrals satisfy a relation analogous to the dilatation equation with scaling degree [2]
@@
From this analysis, it is straightforward to relate the form factor $A$ to the 3K integrals. Indeed, the dilatation Ward identity (18) tells us that the form factor $A$ can be written as a combination of integrals of the following type
@@
For a more detailed review of the topic, see Appendix A and [2, 3, 33]. If the above condition is satisfied, we need to regularize the integral. This can be done by shifting the parameters of the 3K integrals
@@
In general, the regularization parameters $u$ and $v_{i}$ are arbitrary, though in certain cases there can be specific constraints on them.
@@ 5 The conformal $\langle JJO\rangle_{odd}$ and $\langle TTO\rangle_{odd}$  : a reappraisal
Let us now discuss the $\langle JJO\rangle_{odd}$ and $\langle TTO\rangle_{odd}$ correlators, which have been investigated in [15]. The procedure follows the same approach of the $\langle TJO\rangle_{odd}$ discussed in the previous section.
First, we will briefly review the analysis of the $\langle JJO\rangle_{odd}$ correlator. Then, we will proceed with a re-evaluation of the $\langle TTO\rangle_{odd}$, as the primary and secondary equations in the published version of [15] were not correctly identified. Here, we present the correct version of these equations and solve them accordingly. The previous solution reported in [15] for the $\langle TTO\rangle_{odd}$ with $\Delta_{3}=4$ remains valid, and the conclusions and interpretations of that paper are unaffected. However, these corrections have led us to discover new non-vanishing solutions for the $\langle TTO\rangle_{odd}$ characterized by different values of $\Delta_{3}$. Lastly, we further elaborate on all the solutions of both the $\langle JJO\rangle_{odd}$ and $\langle TTO\rangle_{odd}$, offering a new interpretation using the shadow transform.
@@ 5.2 The $\langle TTO\rangle_{odd}$
The most general solution to these equations can then be written in terms of the following combinations of 3K integrals
@@ 5.3 Connections with the anomalies
in the flat limit. By Fourier transforming these expressions, one obtains exactly the solutions in eq. (48) and (5.2) with $\Delta_{3}=4$. Note that $O=\nabla\cdot J_{A}$ is a primary operator since, by acting on it with the special conformal operator $\mathcal{K}$, we obtain a vanishing result
@@
Besides $O=\nabla\cdot J_{A}$, one could alternatively consider $O=g_{\mu\nu}T^{\mu\nu}$ as a scalar operator with $\Delta_{3}=4$. The correlators would then be similarly protected by the possible existence of parity-odd trace anomalies in CFT
@@ 5.4 The shadow transforms and the $\Delta_{3}\leq 0$ solutions
The general solution of these correlators with $\Delta_{3}=4$ and $\Delta_{3}=0$ differ by a factor $p_{3}^{4}$ Although the condition $\Delta_{3}=0$ is non-physical and violate unitarity, we can now ask ourselves if there are any justifications for the existence of such non-vanishing solutions in CFT. As we will see, the conformal and chiral anomalies can still be responsible for such solutions. To demonstrate this, we need to introduce the concept of shadow transform, which we will briefly review [35, 36, 37].
@@
Given this definition, we now re-examine the solutions of the $\langle JJO\rangle_{odd}$ and $\langle TTO\rangle_{odd}$ in eq. (48) and (5.2). For both correlators, the solution with $\Delta_{3}=0$ is connected to the one with $\Delta_{3}=4$ through a shadow transformation of the scalar operator $O$.